OpenStudy (anonymous):

Solve the compound inequality 9 + 3n greater than or equal to 6 or 5n < 25. n less than or greater to −1 or n < 5 n greater than or equal to −5 or n > 5 n greater than or equal to −1 or n < 5 n less than or greater to −5 or n > 5

OpenStudy (anonymous):

9 + 3n >= 6 9 + 3n - 9 >= 6 - 9 3n >= -3 now u can solve this..

OpenStudy (anonymous):

5n < 25 divide both sides by 5

OpenStudy (anonymous):

n<5

OpenStudy (anonymous):

yaaa..

OpenStudy (anonymous):

for the first one..???

OpenStudy (anonymous):

3+n>2

OpenStudy (anonymous):

3+n >= 2 your write substract 3 on both the sides to get n value..

OpenStudy (anonymous):

n>=-1

OpenStudy (anonymous):

yaa ypur write....:)

OpenStudy (anonymous):

next step

OpenStudy (anonymous):

thats all .... now check your option to select the answer.

OpenStudy (anonymous):

Is the answer C.

OpenStudy (anonymous):

yaaa...